# NormXLogit: The Head-on-Top Never Lies

### v1

@@ 4.2 Results
5 Experiment 2: Evidence Alignment
5.1 Experimental Setup
In this section, we focus on per-layer interpretations provided by different methods rather than their global attributions. Specifically, we concentrate on the masked language modeling objective and look for tokens that have the highest impact on the predicted token.
Data.
To assess the target model’s sensitivity to linguistic phenomena, we employ the BLIMP dataset (Warstadt et al. 2020) which contains sentence pairs with minimal contrasts in syntax, morphology, or semantics. The dataset is constructed to provide samples where the true label is uniquely determined by a single word in each sentence. This word, which serves as the decisive factor in determining grammatical acceptability, is termed the evidence. Following Mohebbi et al. (2023), we utilize a subset of the BLIMP dataset, comprising 5 paradigms that represent 3 distinct linguistic phenomena. An example of each phenomenon has been provided in Table 2.
Using spaCy (Honnibal and Montani 2017), we are able to identify the evidence of each linguistic phenomenon. For anaphor number agreement, we employ NeuralCoref777https://github.com/huggingface/neuralcoref to detect the coreferent of the target word. To address determiner-noun agreement, we generate the dependency tree for each sample and extract the determiners corresponding to the target noun. Lastly, for subject-verb agreement, the same dependency tree can be used to identify the subjects associated with the verb.
@@ 5.1 Experimental Setup
In this section, we employ the RoBERTa (Liu et al. 2019) model for our evaluations. We use both pre-trained888The results of the pre-trained model are covered in the Appendix. and fine-tuned versions of the model. For fine-tuning, the target token is replaced with [MASK], and the model is optimized to select the correct target token (the grammatically appropriate word) over the foil token (a similar but grammatically incorrect alternative). Next, for inference, we use the head-on-top, also known as the unembedding matrix, to generate probabilities over the vocabulary.
@@
GlobEnc, ALTI, and Value Zeroing are the attribution methods we consider for comparison in this experiment. Unlike Value Zeroing, which focuses on layer-wise attributions, the other two methods generate global importance scores. To acquire per-layer explanations for GlobEnc and ALTI, we bypass the rollout aggregation method to directly derive per-layer scores and we denote them as GlobEnc¬ and ALTI¬. Moreover, for NormXLogit, to obtain the attributions of layer we use the LogAt on the output representations of -th layer in combination with the norm of the input embeddings where . In this language modeling setup, applying the head-on-top method to the final representations of tokens in the sequence provides a probability distribution over the vocabulary for each input token. Then, is the probability assigned to the target token for each input token, which is considered its attribution score. Using the per-label explanations that could be obtained via LogAt, we also demonstrate the importance of evidence words in predicting both foil and target tokens. These explanations can be generated for any word in the vocabulary, as illustrated in the Appendix. We also consider a random baseline in which tokens attributed equal scores.
@@
Following Yin and Neubig (2022), we define the known evidence as a binary vector with a length equal to the input sequence . In this vector, a value of 1 at a given index indicates the presence of known evidence, while a value of 0 indicates its absence. Similarly, the explanation is also represented as a vector of the same length, where the -th element of it , shows the score produced by an attribution method for predicting the target token. To evaluate the alignment between evidence and explanation vectors, we take advantage of Dot Product and Average Precision metrics.
@@
The dot product computes the total score that the target attribution method assigns to the known evidence.
@@
To evaluate whether an attribution method has found the most important tokens in the input sequence, we use Average Precision. This metric concentrates on the ranking obtained via the attribution method rather than its scores. To calculate the Average Precision for a single sample, we have:
where and indicate the recall and precision at threshold , and is the length of input sequence.
@@ 5.2 Results
As noted earlier, the LogAt scores can be calculated for other tokens in the vocabulary as well. Our analysis shows that the LogAt score for the word ’plural’ () outperforms all other methods in our experiments by a notable margin. This superior performance, unlike that of other random words, might be attributed to the number agreement phenomena underlying this experiment. At layer 7, the results of for Dot Product and Average Precision are and , respectively (refer to Appendix for complete results).
@@ A.1 Computing Infrastructure
All experiments were conducted on a machine with an Nvidia Quadro RTX 8000 GPU with 48GB of memory. The operating system used is Ubuntu 22.04.3 LTS. The ‘requirements.txt’ file included in the code appendix lists all the necessary software libraries and frameworks, along with their respective versions.
@@ A.2 Experiment 1: Complete Results
Figures 6 to 12 illustrate the AOPC and Accuracy across different models and datasets. In Figure 7, we demonstrate the global performance of Value Zeroing on the SST-2 dataset. The results show that this method is not faithful to the model’s decision-making process. This issue may stem from the inherent limitations of the rollout aggregation method, as previously discussed. Additionally, since Value Zeroing is a perturbation-based method, it may also inherit some of the challenges associated with these approaches. For instance, this method zeros out each token’s value vector one at a time, which can lead to problems like ignoring dependencies between features. Consider the following example:
“The movie is mediocre, maybe even bad.”
In this case, erasing “mediocre” or “bad” independently may not significantly impact the overall sentiment of the sentence.
@@ A.3 Experiment 2: RoBERTa Complete Results
The evidence alignment experiment is conducted on a masked language modeling task to understand why a particular target token is chosen. The LogAt method provides per-label attribution scores, enabling us to apply it to other labels (i.e., tokens in the vocabulary) to identify the most important tokens in the input sequence for predicting each specific label. Figures 13 and 14 display the results of the Dot Product and Average Precision alignment metrics for the pre-trained RoBERTa model. An important observation is the notable performance of , which demonstrates its effectiveness in identifying evidence words. This level of performance is not seen with two other randomly chosen words. Specifically, the results are more pronounced in the top layers, indicating that increased context mixing enhances the connection between the evidence and the word “plural.” In other words, as we progress through the layers, the contextualized representation of the evidence word becomes increasingly similar to the word “plural,” resulting in a higher attribution for this word. We attribute the superior performance for the word “plural” primarily to the nature of the phenomena used from the BLIMP dataset, which focused on number agreement. Figures 15 and 16 demonstrate the results for the fine-tuned RoBERTa model.
Appendix B Reproducibility Checklist
This paper:
Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes/partial/no/NA)
Yes. Section 3. Proposed Approach. A detailed conceptual explanation is provided. Also, you can refer to Figure 1(b) which can be helpful to better understand the approach.
Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes/no)
Yes.
Provides well marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper (yes/no)
Yes. We have provided references to fundamental papers and methods where necessary, especially in sections 1. Introduction and 2. Related Work.
Does this paper make theoretical contributions? (yes/no)
No.
Does this paper rely on one or more datasets? (yes/no)
Yes.
A motivation is given for why the experiments are conducted on the selected datasets (yes/partial/no/NA)
Yes. The motivation for the choice of datasets is provided in the ”Data” paragraphs within Sections 4.1 and 5.1 under ”Experimental Setup”.
All novel datasets introduced in this paper are included in a data appendix. (yes/partial/no/NA)
NA. No novel datasets are introduced in this paper.
All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes/partial/no/NA)
NA.
All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. (yes/no/NA)
Yes.
All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. (yes/partial/no/NA)
Yes.
All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. (yes/partial/no/NA)
NA.
Does this paper include computational experiments? (yes/no)
Yes.
Any code required for pre-processing data is included in the appendix. (yes/partial/no).
Yes. The code is provided as supplementary material (code appendix).
All source code required for conducting and analyzing the experiments is included in a code appendix. (yes/partial/no)
Yes.
All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes/partial/no)
Yes.
All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (yes/partial/no)
Partial. All source code includes detailed comments explaining the implementation. However, specific references to the paper for each step are not particularly provided. Instead, different files are organized for each setup to reduce ambiguity and facilitate understanding.
If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. (yes/partial/no/NA)
Yes. We have set seeds for all experiments to ensure reproducibility. The seed values are available in the code appendix separately for each file.
This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. (yes/partial/no)
Yes.
This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. (yes/partial/no)
Partial. This paper formally describes all evaluation and alignment metrics used in both experiments. While the motivation for choosing each metric is not explicitly stated, these metrics are widely recognized in the literature and were selected to facilitate a meaningful comparison between our method and existing approaches.
This paper states the number of algorithm runs used to compute each reported result. (yes/no)
No.
Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. (yes/no)
No.
The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). (yes/partial/no)
No.
This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (yes/partial/no/NA)
Partial. Some hyperparameter details are provided in the paper and the technical appendix, with the remaining details available in the code appendix.
This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (yes/partial/no/NA)
NA. The focus of this paper is not on hyperparameter tuning but rather on investigating a fine-tuned model. Therefore, detailed experimentation with different hyperparameter values was not central to our work. However, for hyperparameters related to attribution methods, we have provided relevant details where applicable.

### v2

@@ 5.1 Experimental Setup
Following Mohebbi et al. (2023), we utilize a subset of the BLiMP dataset comprising five paradigms that represent three distinct linguistic phenomena. Examples of each phenomenon are provided in Table 2. Using spaCy (Honnibal and Montani, 2017), we are able to identify the evidence of each linguistic phenomenon. For anaphor number agreement, we employ NeuralCoref999https://github.com/huggingface/neuralcoref to detect the coreferent of the target word. To address determiner-noun agreement, we generate the dependency tree for each sample and extract the determiners corresponding to the target noun. Lastly, for subject-verb agreement, the same dependency tree can be used to identify the subjects associated with the verb.
@@
In this section, we employ the RoBERTa (Liu et al., 2019) model for our evaluations. We use both pre-trained101010The results of the pre-trained model are covered in Section A.3. and fine-tuned versions of the model. For fine-tuning, the target token is replaced with [MASK], and the model is optimized to select the correct target token (the grammatically appropriate word) over the foil token (a similar but grammatically incorrect alternative). Next, for inference, we use the head-on-top, also known as the unembedding matrix, to generate probabilities over the vocabulary.
@@
GlobEnc, ALTI, and Value Zeroing are the attribution methods we consider for comparison in this experiment. Unlike Value Zeroing, which produces layer-wise attributions, the other two methods generate global importance scores. To acquire per-layer attributions for GlobEnc and ALTI, we bypass the rollout aggregation method to directly derive per-layer scores and we denote them as GlobEnc¬ and ALTI¬. We also consider a random baseline in which tokens are attributed equal scores.
Layer-wise Attribution with NormXLogit.
Following the procedure described in Subsection 3.3, we treat the language modeling setup as a classification task, where the number of labels corresponds to the size of the vocabulary. In this setup, applying the head-on-top to the token representations yields a probability distribution over the vocabulary for each input token. For NormXLogit, to obtain the attributions of layer , we apply LogAt to the output representations of the -th layer, combined with the norm of the input embeddings, where . Then, for each input token, is considered its attribution score with respect to the target token. Thanks to the per-label attributions obtained via LogAt, we can also identify the importance of evidence words in predicting both foil and target tokens. These attributions can be generated for any token in the vocabulary (cf. A.3).
@@
Following Yin and Neubig (2022), we define the known evidence (i.e., ground truth token-level rationale) as a binary vector with a length equal to the input sequence . In this vector, a value of at a given index indicates the presence of known evidence, while a value of indicates its absence. Similarly, the explanation vector has the same length, with each element representing the attribution score assigned to the -th token for predicting the target token. To evaluate the alignment between evidence and explanation vectors, we take advantage of the Dot Product and Average Precision metrics (see Section A.4 for a worked example).
@@
The dot product measures the total score that the target attribution method assigns to the evidence tokens.
@@
To evaluate whether an attribution method successfully identifies the most important tokens, we use Average Precision (AP). This metric concentrates on the ranking obtained via the attribution method rather than its raw scores. For a given sample, AP is computed as:
where and indicate the recall and precision at threshold , and is the length of the input sequence.
@@ 5.2 Results
As noted earlier, the LogAt scores can be calculated for other tokens in the vocabulary as well. Our analysis shows that the LogAt score for the word ‘plural’ () outperforms all other methods in our experiments by a notable margin. This superior performance, unlike that of other random words, might be attributed to the number agreement phenomena underlying this experiment. At layer 7, the results of for Dot Product and Average Precision are and , respectively (cf. A.3).
@@ 6 Conclusion and Future Work
Limitations
As with most attempts to interpret deep learning models, our evaluation, as well as those of existing methods, is inherently constrained by the lack of a definitive gold standard for understanding model internals. To mitigate this, we employed independent and complementary evaluation frameworks, including the AOPC metric, to measure the faithfulness of our approach.
Our work primarily targets models with a transformer architecture, with evaluations focused on models applied to text. Nevertheless, we believe our method is broadly applicable to scenarios where inputs are represented numerically and the model generates token-wise representations.
@@ A.1 Computing Infrastructure
All experiments were conducted on a machine with an Nvidia Quadro RTX 8000 GPU with 48GB of memory. The operating system used is Ubuntu 22.04.3 LTS.
@@ A.2 Experiment 1: Complete Results
Figures 6 to 24 illustrate the AOPC and Accuracy across different models and datasets. In Figure 6, we demonstrate the global performance of Value Zeroing on the SST-2 dataset. The results show that this method is not faithful to the model’s decision-making process. This issue may stem from the inherent limitations of the rollout aggregation method, as previously discussed. Additionally, since Value Zeroing is a perturbation-based method, it may also inherit some of the challenges associated with these approaches. For instance, this method zeros out each token’s value vector one at a time, which can lead to problems like ignoring dependencies between features. Consider the following example:
"The movie is mediocre, maybe even bad."
In this case, erasing “mediocre" or “bad" independently may not significantly impact the overall sentiment of the sentence.
@@
Computational Efficiency.
We conducted experiments to compare the computational efficiency of the attribution methods, using the BERT model with a maximum input length of 512 tokens. Timing results, averaged over five runs and reported in seconds, reveal that the runtime for Gradient Norm and GradientInput remains nearly constant across different input lengths, while Integrated Gradients and DecompX show noticeable increases as input length grows. NormXLogit consistently demonstrates the fastest processing times, with minimal sensitivity to input length. In terms of memory efficiency, we identified the maximum batch size each method could handle within 48GB of GPU memory. NormXLogit significantly outperforms others, processing up to 750 samples per batch, whereas Integrated Gradients and DecompX are limited to very small batch sizes (2 and 1, respectively). Consequently, NormXLogit offers both superior speed and memory efficiency, enabling more scalable attribution computations.
| SST-2 (Acc) | MNLI (Acc) | QNLI (Acc) | |||||||
| Random Baseline | 0.793 | 0.826 | 0.839 | 0.563 | 0.622 | 0.578 | 0.733 | 0.770 | 0.792 |
| Gradient Norm | 0.844 | 0.772 | 0.764 | 0.564 | 0.539 | 0.487 | 0.699 | 0.700 | 0.715 |
| Gradient×Input | 0.818 | 0.751 | 0.758 | 0.536 | 0.507 | 0.501 | 0.681 | 0.684 | 0.716 |
| Integrated Gradients | 0.822 | 0.749 | 0.726 | 0.523 | 0.503 | 0.486 | 0.675 | 0.688 | 0.711 |
| DecompX | N/A | N/A | 0.527 | N/A | N/A | 0.388 | N/A | N/A | 0.627 |
| norm | 0.743 | 0.731 | 0.775 | 0.592 | 0.602 | 0.572 | 0.748 | 0.736 | 0.768 |
| LogAt | 0.768 | 0.791 | 0.757 | 0.554 | 0.599 | 0.545 | 0.715 | 0.724 | 0.745 |
| NormXLogit | 0.720 | 0.711 | 0.684 | 0.471 | 0.526 | 0.427 | 0.649 | 0.584 | 0.653 |
| Method | |||||
| Input Length | Gradient Norm | Gradient×Input | Integrated Gradients | DecompX | NormXLogit |
| 40 | 0.77 | 0.80 | 0.82 | 0.39 | 0.36 |
| 120 | 0.77 | 0.81 | 0.96 | 0.38 | 0.36 |
| 360 | 0.76 | 0.80 | 1.47 | 0.99 | 0.38 |
| 512 | 0.78 | 0.80 | 1.79 | 2.36 | 0.36 |
| Method | |||||
| Evaluation Criteria | Gradient Norm | Gradient×Input | Integrated Gradients | DecompX | NormXLogit |
| Effective Batch Size | 100 | 100 | 2 | 1 | 750 |
| Average Time per Instance (s) | 0.0076 | 0.0081 | 1.4104 | 2.3625 | 0.0005 |
@@ A.3 Experiment 2: RoBERTa Complete Results
The evidence alignment experiment is conducted on a masked language modeling task to understand why a particular target token is chosen. The LogAt method provides per-label attribution scores, enabling us to apply it to other labels (i.e., tokens in the vocabulary) to identify the most important tokens in the input sequence for predicting each specific label. Figures 26 and 26 display the results of the Dot Product and Average Precision alignment metrics for the pre-trained RoBERTa model. An important observation is the notable performance of , which demonstrates its effectiveness in identifying evidence words. This level of performance is not seen with two other randomly chosen words. Specifically, the results are more pronounced in the top layers, indicating that increased context mixing enhances the connection between the evidence and the word “plural." In other words, as we progress through the layers, the contextualized representation of the evidence word becomes increasingly similar to the word “plural," resulting in a higher attribution for this word. We attribute the superior performance for the word “plural" primarily to the nature of the phenomena used from the BLiMP dataset, which focused on number agreement. Figures 28 and 28 demonstrate the results for the fine-tuned RoBERTa model.
A.4 Experiment 2: Worked Example for Alignment Metric Computation
To illustrate how the alignment metrics work, we provide an example using the following sentence:
"Karla thinks/think about it"
In this example, “Karla” serves as the evidence that determines the correct verb (“thinks”).
Vector Definitions.
We first define a binary evidence vector to indicate which tokens are considered known evidence (1 for evidence, 0 otherwise). For this example, the input sequence is tokenized as:
| [Karla, thinks, about, it] |
Since only “Karla” is the evidence, we define:
Next, we define the attribution vector , where each element corresponds to the score assigned by the attribution method to token :
Dot Product:
We compute the dot product between the two vectors:
This score reflects the total attribution the method assigns to the known evidence (“Karla”). A higher dot product indicates better alignment with known evidence.
Average Precision (AP):
This metric evaluates how well the attribution method ranks the evidence token(s). We start by sorting the attribution scores in descending order and recording their indices:
This means the token at index 2 (“about”) has the highest score, followed by index 0 (“Karla”), and so on. The known evidence index (based on ) is:
To compute AP, we iterate through the ranked list and monitor when the evidence index is included. Precision is calculated at each step where recall increases (i.e., when a new evidence token is found):
1.
[2] — does not include index 0 → no change in recall
2.
[2, 0] — includes index 0 → recall changes → precision = 1/2 = 0.5
3.
[2, 0, 1] — no change in recall
4.
[2, 0, 1, 3] — no change in recall
So, the final AP is:
Note that if the evidence index were 2 (evidence ranked first), the AP would be . If it were 1 (ranked third), the AP would be approximately . A higher AP indicates better alignment between the attribution method’s ranking and the true evidence.